If the market for bread is described by the following equations Qd=210-0,6P and Qs=-96+1,1P what is the equilibrium quantity for bread?

1
Expert's answer
2021-06-07T11:30:53-0400

If the market for bread is described by the following equations Qd = 210 - 0.6P and Qs = -96 + 1.1P, then in equilibrium Qd = Qs, so:

"210 - 0.6P = -96 + 1.1P,"

1.7P = 306,

P = 180,

"Q = -96 + 1.1\u00d7180 = 102" units.